Find the equation for circle with center (1,3) that passes through (4, -1)
1
Expert's answer
2012-11-13T07:35:59-0500
The equation of the circle is (x-a)^2+(y-b)^2=r^2,
where (a,b) - coordinates of the centre and r - radius of the circle.
Then (4-1)^2+(-1-3)^2=r^2
25=r^2 then r=5
Finally the equation of the given circle is (x-1)^2+(y-3)^2=25